How much of a 50% salt solution should be added to 80L of a 5% salt solution to obtain a 30% salt solution?
Answer by stanbon(75887)

How much of a 50% salt solution should be added to 80L of a 5% salt solution to obtain a 30% salt solution?
----------
Equation:
salt + salt = salt
0.50x + 0.05*80 = 0.30(x+80)
-------
50x + 5*80 = 30x + 30*80
20x = 25*80
x = 100 L (amt. of 50% solution needed)
-------------------
Cheers,
Stan H.
-------------------